# Adjoint motives of modular forms and the Tamagawa number conjecture

### v1

@@ 8.3. Bloch-Kato conjecture
With the notation in §1.1, we further denote by MF the category of filtered $\phi$-modules as defined in [Fo1, 1.2.1]. We denote by ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} (resp. $K_{\lambda}$-MF) the category of ${\mathcal{O}}_{\lambda}$- (resp. $K_{\lambda}$-) modules in the abelian category $\mathcal{M}\mathcal{F}$}} (resp. additive category MF) and by $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ (resp. $\text{$K_{\lambda}$-MF}^{a}$) the full subcategories with filtration restrictions as in §$1.$ Scalar extension -$\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$ induces an exact functor $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}\rightarrow\text{$K_{\lambda}$-MF}$ where the notion of exactness in MF is defined in [Fo1, 1.2.3].
@@
These arguments apply to ${\mathcal{D}}_{1}={\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ which is an object of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{0}$ if $\lambda\in S_{f}$, more specifically if $\ell\nmid N$ and $\ell>k$. We have $T_{1}=T_{2}={\mathcal{M}}_{f,\lambda}$ and $T={\mathcal{A}}_{f,\lambda}\oplus{\mathcal{O}}_{\lambda}$. Our choice of $\omega_{A}$ then ensures that $\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{A}}^{0}({\mathcal{A}}_{f,\lambda})={\mathcal{O}}_{\lambda}$. For $B_{f}=A_{f}(1)$ we can use the same argument as long as both ${\mathcal{D}}_{1}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ and ${\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}[1]$ are objects of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{1}$. This is the case if $\ell>k+1$ or if $l=k+1$ and ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient AA in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
@@
If $a_{\ell}\equiv 0\mod\lambda$ then ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient AA in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
@@
and $\dim_{{\mathcal{O}}_{\lambda}/\lambda}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}=1$ the map $\bar{\phi}$ is nonzero, hence conjugate to $\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)$. Since $\bar{\phi}=l^{k-1}\bar{\phi}^{k-1}=0$ on $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}$ and because of (76) we have $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}=\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$. It is now easy to see that $\bar{{\mathcal{M}}}$ is a simple object in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}}: Any proper subobject $N\subset\bar{{\mathcal{M}}}$ is $\bar{\phi}$-stable, hence contained in $\text{ker}(\bar{\phi})=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}$ and we have $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=N$. But again by (76) we find $\bar{\phi}^{k-1}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}})\not\subseteq\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$ so that $N=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=0$. If AA is a nonzero quotient of $\mathcal{M}$} then $\bar{A}$ is a nonzero quotient of $\bar{{\mathcal{M}}}$ hence equal to $\bar{{\mathcal{M}}}$ and we find $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{A}\neq\bar{A}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A\neq A$. ∎
@@
We shall first prove Theorem 8.10 for $M=B_{f}=A_{f}(1)$ in which case Lemma 8.11 applies. Fix a prime $\lambda\in S_{f}$ and let $\Sigma$ be the set of primes dividing NN if $N>1$ or put $\Sigma=\{p\}$ for some prime $\lambda\nmid p$ if $N=1$. The isomorphism $\gamma:M_{f}\rightarrow M_{f}^{\Sigma}$ of Proposition 6.3 satisfies
@@
where $\eta^{\Sigma}_{f}$ was defined before Proposition $6.$ On the other hand by (78), Theorem 8.5 and the remark after the proof of Theorem 8.5 we have for $\lambda\in S_{f}$

### v2

@@ 8.3. Bloch-Kato conjecture
With the notation in §1.1, we further denote by MF the category of filtered $\phi$-modules as defined in [Fo1, 1.2.1]. We denote by ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} (resp. $K_{\lambda}$-MF) the category of ${\mathcal{O}}_{\lambda}$- (resp. $K_{\lambda}$-) modules in the abelian category $\mathcal{M}\mathcal{F}$}} (resp. additive category MF) and by $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ (resp. $\text{$K_{\lambda}$-MF}^{a}$) the full subcategories with filtration restrictions as in §1.1. Scalar extension -$\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$ induces an exact functor $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}\rightarrow\text{$K_{\lambda}$-MF}$ where the notion of exactness in MF is defined in [Fo1, 1.2.3].
@@
These arguments apply to ${\mathcal{D}}_{1}={\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ which is an object of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{0}$ if $\lambda\in S_{f}$, more specifically if $\ell\nmid N$ and $\ell>k$. We have $T_{1}=T_{2}={\mathcal{M}}_{f,\lambda}$ and $T={\mathcal{A}}_{f,\lambda}\oplus{\mathcal{O}}_{\lambda}$. Our choice of $\omega_{A}$ then ensures that $\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{A}}^{0}({\mathcal{A}}_{f,\lambda})={\mathcal{O}}_{\lambda}$. For $B_{f}=A_{f}(1)$ we can use the same argument as long as both ${\mathcal{D}}_{1}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ and ${\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}[1]$ are objects of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{1}$. This is the case if $\ell>k+1$ or if $l=k+1$ and ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient $A$ in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
@@
If $a_{\ell}\equiv 0\mod\lambda$ then ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient $A$ in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
@@
and $\dim_{{\mathcal{O}}_{\lambda}/\lambda}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}=1$ the map $\bar{\phi}$ is nonzero, hence conjugate to $\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)$. Since $\bar{\phi}=l^{k-1}\bar{\phi}^{k-1}=0$ on $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}$ and because of (76) we have $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}=\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$. It is now easy to see that $\bar{{\mathcal{M}}}$ is a simple object in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}}: Any proper subobject $N\subset\bar{{\mathcal{M}}}$ is $\bar{\phi}$-stable, hence contained in $\text{ker}(\bar{\phi})=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}$ and we have $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=N$. But again by (76) we find $\bar{\phi}^{k-1}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}})\not\subseteq\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$ so that $N=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=0$. If $A$ is a nonzero quotient of $\mathcal{M}$} then $\bar{A}$ is a nonzero quotient of $\bar{{\mathcal{M}}}$ hence equal to $\bar{{\mathcal{M}}}$ and we find $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{A}\neq\bar{A}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A\neq A$. ∎
@@
We shall first prove Theorem 8.10 for $M=B_{f}=A_{f}(1)$ in which case Lemma 8.11 applies. Fix a prime $\lambda\in S_{f}$ and let $\Sigma$ be the set of primes dividing $N$ if $N>1$ or put $\Sigma=\{p\}$ for some prime $\lambda\nmid p$ if $N=1$. The isomorphism $\gamma:M_{f}\rightarrow M_{f}^{\Sigma}$ of Proposition 6.3 satisfies
@@
where $\eta^{\Sigma}_{f}$ was defined before Proposition 6.6. On the other hand by (78), Theorem 8.5 and the remark after the proof of Theorem 8.5 we have for $\lambda\in S_{f}$